### v2

# MlingConf: A Comprehensive Study of Multilingual Confidence Estimation on Large Language Models

Abstract
The tendency of Large Language Models (LLMs) to generate hallucinations and exhibit overconfidence in predictions raises concerns regarding their reliability. Confidence or uncertainty estimations indicating the extent of trustworthiness of a model’s response are essential to developing reliable AI systems. Current research primarily focuses on LLM confidence estimations in English, remaining a void for other widely used languages and impeding the global development of reliable AI applications. This paper introduces a comprehensive investigation of Multilingual Confidence estimation (MlingConf) on LLMs. First, we introduce an elaborated and expert-checked multilingual QA dataset. Subsequently, we delve into the performance of several confidence estimation methods across diverse languages and examine how these confidence scores can enhance LLM performance through self-refinement. Extensive experiments conducted on the multilingual QA dataset demonstrate that confidence estimation results vary in different languages, and the verbalized numerical confidence estimation method exhibits the best performance among most languages over other methods. Finally, the obtained confidence scores can consistently improve performance as self-refinement feedback across various languages.
MlingConf: A Comprehensive Study of Multilingual Confidence Estimation on Large Language Models
Boyang Xuethanks:   Equal Contribution. , Hongru Wang♠∗, Rui Wang, Sheng Wang Zezhong Wang, Yiming Du, Kam-Fai Wong The Chinese University of Hong Kong The University of Hong Kong {byxue, hrwang, kfwong}@se.cuhk.edu.hk
1 Introduction
Although LLMs have exhibited noteworthy proficiencies across a range of tasks, their susceptibility to generating hallucinated contents and being overconfident in their generations incurs concerns about unreliability in real-world applications Ji et al. (2023); Rawte et al. (2023). Therefore, it becomes increasingly crucial for users to directly ascertain how much they can trust a model’s response. Assessing the confidence or uncertainty of a model’s output can immediately indicate the level of reliability to users, thereby playing a key role in developing trustworthy AI systems Geng et al. (2023); Kadavath et al. (2022).
However, existing research on confidence or uncertainty estimations of LLMs is confined exclusively in English Lin et al. (2023, 2022); Chen and Mueller (2023); Duan et al. (2023); Varshney et al. (2023); Tian et al. (2023b). With the accelerated pace of globalization, it is essential to develop worldwide reliable AI applications. There is a dearth of research dedicated to extending the confidence estimation methods of LLMs reported in English to encompass other widely used languages such as Chinese and French. The limitation raises a question: How does the confidence estimation performance vary across various languages? By performing several confidence estimation methods, the optimal approach applicable within multilingual environments can be determined. When viewing different languages, we can discern if different languages have consistent confidence scores on LLMs. In addition, since the uncertainty estimation can be used as feedback for LLM self-refinement to evaluate its further use to improve the the previously proposed answers of LLMs (Yang et al., 2023), we also explore the potential variations in different languages when using the obtained scores for LLM self-refinement Madaan et al. (2023).
To this end, we introduce MlingConf, a comprehensive investigation of multilingual confidence estimation for LLMs. First, we meticulously construct a high-quality benchmark for multilingual confidence estimation, named Mling-TriviaQA, derived from TriviaQA Joshi et al. (2017). The dataset encompasses ten diverse languages and is rigorously checked by linguistic experts. We employ several confidence estimation methods on LLMs including likelihood-based confidence Varshney et al. (2023), true probability confidence Kadavath et al. (2022), and linguistic confidence Lin et al. (2022); Xiong et al. (2023) across a spectrum of languages. Then we treat the obtained confidence scores as feedback of self-refinement experiments to observe performance improvements across diverse languages. We demonstrated three principal findings across our experiments: 1) Confidence estimation results vary in different languages; 2) The verbalized numerical confidence estimation method consistently outperforms other methods across most languages; 3) The derived confidence scores reliably enhance performance when utilized as self-refinement feedback in diverse linguistic contexts.
The contributions are summarized as follows:
$\bullet$ To the best of our knowledge, this study is the first to investigate multilingual confidence estimations on LLMs. We introduce an intricately constructed and expert-checked multilingual QA dataset, denominated Mling-TriviaQA, to serve as the multilingual confidence estimation benchmark.
$\bullet$ We present a comprehensive study on various languages with several confidence estimation methods, as well as employ confidence scores as self-refinement feedback. Several findings are presented which provide a valuable insight for future reliable multilingual research.
2 Related Works
Both confidence and uncertainty indicate the level of certainty or assurance of responses generated by LLMs and are occasionally regarded interchangeably Geng et al. (2023). Numerous methodologies employing logit-based uncertainty estimation techniques on white-box LLMs have been extensively utilized in Tian et al. (2023a); Vazhentsev et al. (2023). Kadavath et al. (2022) instructs the LLM to self-evaluate the correctness of the generated answer by directly accessing $p(\mathrm{True})$. Kuhn et al. (2023) introduces the concept of semantic entropy, aggregating invariant token sequences by shared meanings. Recent works pay particular attention to linguistic confidence via prompting LLMs to express certainty in verbalized numbers or words for the black-box LLMs. Lin et al. (2022) instructs LLMs to produce a level of confidence in natural language. Xiong et al. (2023) focuses on confidence elicitation on LLMs to accurately express confidence.
Most recent LLMs primarily pre-trained on English corpora have showcased remarkable capabilities Pires et al. (2019); Shi et al. (2023); OpenAI (2023); Workshop et al. (2023); Chowdhery et al. (2022); Zhang et al. (2022); Brown et al. (2020). However, their efficacy in other languages remains limited, hindering the global proliferation of LLM deployment and provoking a surge in research interest. The previous works only consider the confidence estimation in English. In this work, we systematically investigate multilingual confidence estimations on both black-box and white-box LLMs.
3 Confidence Estimation Methods
In this section, we investigate several commonly used confidence estimation methods for generative LLMs. Specifically, we denote $\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})$ as the confidence score associated with the output sequence $\boldsymbol{y}=[y_{1},y_{2},\dots,y_{N}]$ given the input context $\boldsymbol{x}=[x_{1},x_{2},\dots,x_{M}]$.
Likelihood-based Confidence:
The likelihood-based confidence is estimated by calculating the joint token-level probabilities over $\boldsymbol{y}$ conditioned on $\boldsymbol{x}$. As longer sequences are supposed to have lower joint likelihood probabilities that shrink exponentially with length, we calculate the geometric mean by normalizing (Likelihood-Norm), arithmetically averaging (Likelihood-Avg) the product of conditional token probabilities in the output by the sequence length. We also employ the minimum of token probabilities. (Likelihood-Min) Their respective confidence can be represented as:
$\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})$ $=\left({\prod^{N}_{i}p(y_{i}|\boldsymbol{y}_{<i},\boldsymbol{x})}\right)^{\frac{1}{N}}$
$\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})$ $=\frac{1}{M}{\sum^{M}_{i}p(y_{i}|\boldsymbol{y}_{<i},\boldsymbol{x})}$ (1)
$\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})$ $={\min\left\{p(y_{1}|\boldsymbol{x}),\dots,p(y_{N}|\boldsymbol{y}_{<N},\boldsymbol{x})\right\}}$
True Probability Confidence:
The True Probability confidence score is implemented by simply asking the model itself if its first proposed answer $\boldsymbol{y}$ to the question $\boldsymbol{x}$ is true Kadavath et al. (2022), and then obtaining the probability $p(\mathrm{True})$, which can implicitly reflect self-reflected certainty.
Self-verbalized Confidence:
We adopt verbalized numerical probability (Verbal Number) and words (Verbal Word) in token-level space as LLM’s confidence estimations. The verbalized word contains a set of five words (“lowest”, “low”, “medium”, “high”, “highest”) indicating the confidence degrees Lin et al. (2022); Xiong et al. (2023).
An illustration of the above confidence estimation methods are presented in Figure 5 in Appendix.
4 Dataset Construction
Owing to the lack of multilingual resources to comprehensively exhibit confidence estimation across diverse languages, our study leverages an extensive multilingual dataset, namely Mling-TriviaQA derived from TriviaQA (English) (Joshi et al., 2017) across ten languages, encompassing English, Japanese, Chinese, Korean, French, Arabic, German, Indonesian, Thai and Italian on which we have employed linguistic experts to check the translation quality. These languages are selected according to the language family category as in Appendix B. We extract and translate numerous closed-book question-answer pairs from TriviaQA which is widely used in confidence estimation Kuhn et al. (2023); Lin et al. (2023). The language family and ISO code information of the investigated languages and why we employ these languages for research are listed in Appendix B.
Figure 1: Data collection of Mling-TriviaQA.
Lang. Accu. Lang. Accu. Lang. Accu.
it 98% ko 96% ja 96%
fr 100% zh 100% de 100%
ar 98% id 98% th 96%
Table 1: Translation accuracy evaluated by linguistic experts on 50 randomly selected samples.
The construction of the Mling-TriviaQA dataset in this study follows a three-stage procedure as delineated in Figure 1. ➀: the QA examples extracted from the TriviaQA dev set are translated into different languages through both GPT-4 OpenAI (2023)  [^1] https://platform.openai.com/docs/api-reference and DeepL  [^2] https://www.deepl.com/pro-api; ➁: we gauge the similarity scores of the results by two translators using the uni-gram F1-measure, retaining only the top 2k samples for further experiments; ➂: after filtering, we employ several experts majoring in linguistics to examine the translation performance across 50 randomly selected samples as shown in Table 1. Human evaluation results suggest the translations are high-quality for experiments. The dataset details are presented in Appendix A.
5 Experiments
5.1 Evaluation Metrics
Following Filos et al. (2019), we employ Area Under the Receiver Operator Characteristic Curve (AUROC) to assess the effectiveness of confidence estimation by quantifying how likely a correct answer possesses a higher uncertainty score than an incorrect one, implement by sklearn toolkit  [^3] https://github.com/scikit-learn/scikit-learn/blob/main/sklearn/metrics/_ranking.py.
For evaluating the answer quality of LLMs, we compute the F1 score, exact matching (EM), and the NLI score as accuracy between the generated answer and gold reference, where the NLI score (Correct scores 1 and Incorrect scores 0) is assessed using GPT-3.5 Turbo and the multilingual prompts are provided in Appendix C.
5.2 Implementation Details
The experiments are carried out using three LLMs: GPT-3.5 Turbo, GPT-4 OpenAI (2023) and PaLM 540B Chowdhery et al. (2022). Few-shot prompts are utilized for answer generation, with a fixed temperature setting of $T=0.8$.
5.3 Results and Analysis
Languages en zh fr ko de ja it ar id th Avg
GPT-3.5 Turbo
Avg Length 4.71 15.87 9.22 15.51 8.03 12.96 10.58 14.79 7.75 22.72 -
Likelihood-Min 78.53 83.66 72.33 80.90 79.97 81.30 74.60 82.36 78.37 82.30 79.43
Likelihood-Avg 77.84 83.77 71.65 83.09 77.70 80.77 71.95 80.26 76.51 76.89 78.04
Likelihood-Norm 78.01 84.59 72.14 83.21 78.58 81.32 73.08 81.51 77.37 79.36 79.11
True Probability 83.97 94.55 80.59 93.20 81.49 92.26 80.77 93.46 86.08 92.90 87.92
Verbal Number 87.14 94.69 83.47 92.91 87.26 92.18 84.94 92.09 88.64 89.33 89.24
Verbal Word 85.60 89.04 82.48 87.08 85.05 88.63 81.48 87.61 86.15 86.39 85.95
GPT-4
Avg Length 4.30 7.90 5.95 8.61 5.62 7.72 5.47 14.79 5.25 11.15 -
Likelihood-Min 67.40 77.80 73.63 76.38 70.04 75.85 69.28 78.35 71.90 75.57 73.59
Likelihood-Avg 68.08 76.78 74.75 76.92 70.92 76.45 69.78 79.29 72.40 77.61 74.29
Likelihood-Norm 67.88 76.70 74.32 76.83 70.91 76.35 69.67 79.10 72.36 77.09 74.11
True Probability 87.33 91.23 84.03 92.60 87.44 95.37 85.44 89.13 82.66 91.29 88.65
Verbal Number 89.97 96.13 85.03 96.14 89.89 94.92 88.85 92.65 87.19 91.13 91.18
Verbal Word 82.24 84.48 82.41 89.04 85.45 92.59 80.83 91.36 86.65 90.25 86.53
PaLM 540B
Avg Length 5.18 12.44 7.32 11.05 6.93 10.86 10.34 15.22 5.84 18.05 -
Verbal Number 82.31 86.65 80.44 87.49 83.42 88.60 81.77 86.50 79.37 86.12 84.26
Verbal Word 79.36 83.60 77.44 81.26 78.45 82.19 77.50 78.36 78.45 80.69 79.73
Table 2: Experimental results of AUROC $\uparrow$ on Mling-TriviaQA set using several confidence estimation methods. Likelihood-based cannot be conducted on PaLM 540B since it is black-boxed and unable to provide probabilities.
We present the results of the confidence estimation performance across different languages. The key findings are listed below.
a. The confidence estimation results vary in different languages.
This phenomenon can be attributed to several factors. Despite the powerful multilingual capacity of state-of-the-art LLMs, discrepancies persist in the quantity of distinct linguistic training corpora available for each language. Furthermore, we report a correlation between confidence estimation performance and the lengths of generated sequences across various languages. Notably, languages characterized by longer generated sequences, such as Chinese and Thai, demonstrate superior confidence estimation performance compared to those belonging to the Indo-European language family, such as German. This trend can be attributed to the notion that an increased number of tokens inherently contains more information, thereby aiding in the reduction of uncertainties.
b. The verbalized numerical confidence estimation method emerges as the optimal method among all other estimation methods.
As shown in Table 2, self-verbalized numerical confidence attains prominence across most languages, yielding the highest AUROC scores on three LLMs. Previous studies only validate the efficacy of self-verbalized confidence assessment exclusively within English. Our findings indicate that the Verbal Number approach is also preferable in other languages. Consequently, we have opted to employ this approach for subsequent experiments.
c. Confidence scores as self-refinement feedback can consistently improve performance across various languages.
As shown in Figure 2, we apply random sampling feedback based on confidence scores in all investigated languages. The confidence scores acquired through this process serve as a fine-grained quantified indicator, guiding the LLM to correct or optimize its answers. We randomly sample a float $k$ from a uniform distribution $K\sim\mathrm{U}(0,1)$ as the threshold. The self-refinement process will be activated when $\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})\leq k$ (threshold selection strategies are reported in Appendix D). The results underscore the efficacy of employing confidence scores as feedback for self-refinement to improve performance across all languages using two LLMs. Notably, languages with higher-quality confidence scores such as Chinese and Japanese, exhibit substantial improvements in 12.50% and 11.25% absolute accuracy gains. The evaluations of EM and F1 for self-refinement are presented in Appendix D.
Figure 2: Accuracy of responses generated by (a) GPT-3.5 Turbo and (b) GPT-4 before and after self-refinement with confidence scores as feedback.
6 Conclusion
This paper presents a comprehensive investigation of multilingual confidence estimation MlingConf on LLMs, encompassing the construction of the self-curated Mling-TriviaQA dataset and experiments with various confidence estimation methods in multiple languages. Several valuable findings elucidate the performance of confidence estimation across diverse languages. By addressing the limitations of existing confidence estimation research confined to English, this study makes a significant contribution towards the advancement of confidence estimation techniques to generalize across diverse linguistic contexts, thereby enhancing the reliability of the global AI system.
Limitations and Future Work
The limitations and prospects for future research are outlined as follows:
1. The present study is constrained by the substantial cost associated with the API cost using GPT-4 and DeepL for translation as well as human verification checks by linguists. In order to ensure the data quality, our exploration into multilingual research is restricted to ten languages in the first version of the benchmark. This initial phase aims to delve into confidence estimation within multilingual domains. Our future endeavors will involve the expansion of the benchmark dataset, encompassing additional languages and data points to enrich our investigations.
2. Our primary experiments are confined to the utilization of widely employed models GPT-3.5, GPT-4, and PaLM 2. Moreover, certain white-box LLMs, such as LLaMA-2 and BLOOM underperform in some low-resource languages like Thai due to inadequate or absent multilingual training corpora. Moving forward, we will broaden the scope of our research by incorporating additional available LLMs, thereby facilitating a more comprehensive analysis of confidence estimation methodologies.
Ethics Statement
In this paper, we introduce a self-constructed multilingual QA dataset derived from the publicly available TriviaQA dataset named Mling-TriviaQA. The selection of investigated languages in this work depends on whether we can employ appropriate linguistic experts. Most linguistic specialists are M.Phil. or Ph.D. students majoring in linguistics and others are from crowd-sourcing platforms. We meticulously adhered to legal and ethical standards throughout the data collection process, prioritizing privacy and obtaining informed consent. Linguistic experts were furnished with comprehensive details regarding the study’s objectives, data collection methodologies, and associated risks or benefits. They were afforded the opportunity to seek clarifications and voluntarily provide consent before their involvement. All collected data were exclusively utilized for research purposes.
Appendix A Dataset Details
A.1 TriviaQA Dataset
The TriviaQA dataset Joshi et al. (2017) is a realistic text-based reading comprehension question-answering dataset containing 650K question-answer-evidence triples from 95K documents collected from Wikipedia and the websites, served as a benchmark for evaluating machine comprehension and question-answering systems, which is more challenging than standard QA benchmark datasets where the answer spans can be directly retrieved and copied.
Appendix B Language Information
The basic information including the language family, ISO code, Macroarea of the investigated languages is presented in Table 3. The languages investigated in this work are selected following two principles.
  1. 1.
    For each selected language, we can employ one linguistic expert for the human check to ensure the data quality.
  2. 2.
    The candidate languages are selected based on language family. A language family represents a collective of cognate languages stemming from a common ancestral source, serving as a focal point within the domain of linguistics  [^4] https://en.wikipedia.org/wiki/Language_family. Some well-known language families include Indo-European and Afro-Asiatic. Linguists posit that languages within a shared language family exhibit congruent patterns of cultural cognition and thought processes Bouckaert et al. (2012); Calude and Pagel (2011). Motivated by the premise, we develop three distinct categorizations based on language family affiliations to determine the optimal strategy for cross-lingual confidence estimations. As English belongs to Indo-European languages, we introduce the other four Indo-European languages including Italian, Indonesian, French, German, and five non-Indo-European languages including Chinese, Japanese, Arabic, Korean and Thai.
Language ISO 639-3 ISO 639-1 Family Macroarea
English eng en Indo-European Eurasia
Italian ind it Indo-European Eurasia
French fra fr Indo-European Eurasia
German deu de Indo-European Eurasia
Indonesian ind id Austronesian Eurasia
Chinese zhs zh Sino-Tibetan Eurasia
Japenese jpn ja Japanese-Ryukyuan Eurasia
Arabic arb ar Afro-Asiatic Eurasia
Korean kor ko Koreanic Eurasia
Thai tha th Kra–Dai Eurasia
Table 3: List of International Standard Organization (ISO) 639-1 Codes.
Appendix C Prompt Details
Figure 3: Prompt templates for answering questions for each language.
Figure 4: Prompt templates for Self-aware confidence estimation for each language.
Multilingual prompts for answer generation and self-estimated confidence estimation are presented in Figure 3 and 4. Notably, the prompts for self-reflected true probability confidence estimation are followed by previous work Kadavath et al. (2022); Kuhn et al. (2023).
Appendix D Self-refinement Experiments
Experimental results of EM and F1 scores between gold answers and responses generated by (a) GPT-3.5 Turbo before and after self-refinement with confidence scores as feedback on Mling-TriviaQA dataset are presented in Figure 6. The trends are similar to the accuracy improvements in Figure 2.
Figure 5: An illustration of several confidence estimation methods on LLMs. $\star$ Note that all complete multilingual prompts used in this work are presented in Appendix C.
Lang. Feedback EM $\uparrow$ F1 $\uparrow$ Accu. $\uparrow$
$\texttt{Conf}\leq k$
en First proposed 63.60 73.12 72.30
$k=0.25$ 63.25 72.85 73.50
$k=0.50$ 64.50 74.18 74.35
$k=0.75$ 64.85 74.12 74.65
$k=1.00$ 35.25 49.32 60.20
$k\sim\mathrm{\mathrm{U}(0,1)}$ 64.60 74.46 74.50
zh First proposed 16.90 17.14 23.55
$k=0.25$ 16.95 17.32 27.60
$k=0.50$ 17.25 19.44 32.75
$k=0.75$ 17.55 21.00 34.85
$k=1.00$ 11.70 12.82 14.30
$k\sim\mathrm{\mathrm{U}(0,1)}$ 18.40 21.36 38.70
fr First proposed 31.50 47.19 47.20
$k=0.25$ 31.60 47.06 48.50
$k=0.50$ 32.45 47.68 50.35
$k=0.75$ 32.85 48.15 51.65
$k=1.00$ 19.80 24.37 30.20
$k\sim\mathrm{\mathrm{U}(0,1)}$ 32.80 48.24 53.85
Table 4: Experimental results of self-refinement with different confidence estimation feedback on English, Chinese, and French set on GPT-3.5 Turbo.
D.1 Threshold Selection for Self-Refinement
Feedback
We have further applied the confidence scores obtained as fine-grained feedback signals for self-refinements. We employ two methods to determine the feedback utilized: fixed-value and random sampling approaches. The fixed-value method introduces a threshold $k$, and the answers first proposed by the LLM with confidence score $\texttt{Conf}(\boldsymbol{x},\boldsymbol{y})\leq k$ will be self-refined by the LLMs. We conduct experiments with four thresholds spanning {0.25, 0.50, 0.75, 1.00}. In contrast, random sampling determines feedback by randomly sampling a value $k$ from a uniform distribution $k\sim\mathrm{U}(0,1)$. Similarly, the self-refinement process is activated for LLM’s first proposed answers with confidence scores $\texttt{Conf}(\boldsymbol{x},\boldsymbol{y})\leq k$.
Random sampling outperforms fixed-value threshold as confidence score feedback for LLM self-refinement.
As shown in Table 4, we have investigated two methods to determine feedback in three languages encompassing English, Chinese, and French. For fixed threshold, self-refining the outputs of those with self-verbalized numerical confidence scores less than 0.75 yields significant performance enhancements. When the threshold is set to 1.00 which mandates self-refinement for all generated outputs, the performance dramatically decreases as many correct answers are inadvertently modified. Better or comparable performance is obtained using random sampling over the fixed-value method in all tested languages. Therefore, we have elected to employ the random sampling method as confidence estimation feedback for further self-refinement experiments in other languages.
Figure 6: EM and F1 scores of different languages before and after self-refinement using confidence scores as feedback.